### v1

@@ Abstract
Humans generally acquire new skills without compromising the old; however, the opposite holds for Large Language Models (LLMs), e.g., from LLaMA to CodeLLaMA. To this end, we propose a new post-pretraining method for LLMs with an expansion of Transformer blocks. We tune the expanded blocks using only new corpus, efficiently and effectively improving the model’s knowledge without catastrophic forgetting. In this paper, we experiment on the corpus of code and math, yielding LLaMA Pro-8.3B, a versatile foundation model initialized from LLaMA2-7B, excelling in general tasks, programming, and mathematics. LLaMA Pro and its instruction-following counterpart (LLaMA Pro - Instruct) achieve advanced performance among various benchmarks, demonstrating superiority over existing open models in the LLaMA family and the immense potential of reasoning and addressing diverse tasks as an intelligent agent. Our findings provide valuable insights into integrating natural and programming languages, laying a solid foundation for developing advanced language agents that operate effectively in various environments.
[Uncaptioned image] LLaMA Pro: Progressive LLaMA with Block Expansion
[…]
@@ 2 Related Work
The field of large language models has witnessed significant progress in recent years. The growth in model and data scale has played a crucial role in achieving state-of-the-art performance across various tasks Hoffmann et al. (2022); Kaplan et al. (2020); Chowdhery et al. (2023). Concurrently, the development of more generalist models has led to the creation of models that can address diverse problems and quickly adapt to new tasks Radford et al. (2019); Brown et al. (2020). These advancements have been further bolstered by the open-source community, which has released powerful open large language models for research, such as LLaMA Touvron et al. (2023) and CodeLLaMA Roziere et al. (2023). Our work builds upon these developments by providing a methodology for specializing large language models in the domain of code, paving the way for future research and applications in this area.
@@
Language model applications typically involve a two-step process: an initial general-domain pretraining step, followed by domain-specific training Roziere et al. (2023); Azerbayev et al. (2023). The fine-tuning step is often aimed at enhancing instruction-following abilities Sanh et al. (2021); Wei et al. (2021); Wang et al. (2023d) or aligning the model’s outputs with human preferences Ziegler et al. (2019); Ouyang et al. (2022); Bai et al. (2022). Additionally, some studies explore adapting pretrained models to novel domains using parameter-efficient fine-tuning methods Houlsby et al. (2019); Hu et al. (2021); Wu et al. (2023a). Many works also focus on how to do continual learning after the pretraining phace Wang et al. (2023b); Gupta et al. (2023); Scialom et al. (2022). In our work, we propose an adaptation strategy that combines continued training with targeted general capability maintenance, allowing large language models to specialize in specific tasks without sacrificing their overall performance.
@@
In recent years, progressive training has gained attention for its ability to accelerate the training of large-scale models in both computer vision Zhang et al. (2023) and NLP research Yao et al. (2023); Li et al. (2023b).  Gong et al. (2019) proposed a stacking method that doubles the model depth at each stage. CompoundGrow Gu et al. (2020) extends stacking by incorporating FeedForward Network (FFN) expansion into the schedule design.  Shen et al. (2022) proposed a staged method that further supports expanding the hidden size of features. Bert2BERT Chen et al. (2021a) and LiGO Wang et al. (2023a) support all possible growth dimensions. Our method employs depth growth to preserve general performance while adapting to a specific domain.
@@ 3.1 Preliminaries: The LLaMA Block
The FFN block in the LLaMA model utilizes the SwiGLU activation function, which can be defined as:
@@ 3.2 Block Expansion
Given a model with blocks $(\phi_{0},\phi_{1},...,\phi_{L})$, the block expansion incorporates an identity block $\phi_{id}$ after each block in the original model, ensuring that the expanded model maintains the same output after expansion. The identity block is defined as $\phi_{id}(x)=x$, where the input and output are identical.
@@
The entire training pipeline is depicted in Figure 2. Our method concentrates on the post-pretraining stage, targeting specific domain corpora such as code corpora. We begin by initializing our model with large language models trained on extensive unlabeled general corpora, where all blocks will be fine-tuned. To enhance the model’s capacity for accommodating additional domain knowledge while retaining its general knowledge, we employ block expansion to increase the number of blocks in the LLM. During this process, we only fine-tune the newly added blocks while freezing the original blocks, thereby preserving the general abilities of the model.
@@ 4 Experiments
This section presents our key experimental findings. We begin with experimental settings (described in Sec. 4.1), and then verify the effectiveness of block expanded tuning after pretraining (described in Sec. 4.2). Next, we give the supervised finetuning (SFT) results (described in Sec. 5). Finally, ablation studies of the key design choices are presented (described in Sec. 6).
@@ 4.1 Experimental Settings
We construct a dataset that concentrates on code and math. For the code component, we rely on the Stack-dedup dataset, which is a compilation of permissively licensed source codes from GitHub. Among all the programming languages available in Stack-dedup, we specifically utilize the Python split. As for the math component, we opt for the Proof-pile-2 dataset Azerbayev et al. (2023), a 55-billion-token amalgamation of scientific papers, web data containing mathematical content, and mathematical code.
Data source Tokens Weight
Proof-Pile-2 55B 1.00
 AlgebraicStack 11B
 OpenWebMath 15B
 ArXiv 29B
The-Stack-Dedup
 Python 22B 1.50
@@
During the instruction fine-tuning phase, we combine five data sources to create LLaMA Pro - Instruct. These sources include ShareGPT [^1] https://huggingface.co/datasets/anon8231489123/ShareGPT_Vicuna_unfiltered, which contains real user and ChatGPT chat history records, and the WizardLM evolution instruction dataset Xu et al. (2023a), offering a wealth of instruction data with varying complexity levels. We also incorporate the evolution CodeAlpaca dataset Luo et al. (2023), which includes complex coding tasks generated by ChatGPT and their corresponding solutions. Additionally, we use MetaMath Yu et al. (2023), which reframes questions from multiple perspectives, and SlimOrca Lian et al. (2023), a curated subset of our OpenOrca data. SlimOrca provides an efficient route to achieve performance comparable to using larger data slices, while only incorporating approximately 500,000 GPT-4 completions.
ShareGPT User prompts GPT-3.5/GPT-4 63,817 2.9 293.2 1157.1
WizardLM_evol_instruct_V2 GPT-4 GPT-4 143,000 1.0 602.6 1704.9
SlimOrca Human-written GPT-4 517,982 1.0 574.3 599.3
MetaMath Human-written/GPT-4 GPT-4 395,000 1.0 209.4 498.2
Evol-CodeAlpaca GPT-4 GPT-4 111,272 1.0 652.5 1552.0
The final sft dataset consists of approximately 1M samples. To fine-tune the basic models, we employ specific configurations, including a batch size of 128, a sequence length of 4096, 0.03 warmup ratio, a learning rate of 2e-5, a Cosine learning rate scheduler, and bf16 mixed precision.
@@
We conduct a comparative analysis of LLaMA Pro with the latest state-of-the-art (SOTA) Large Language Models (LLMs). The evaluation is performed on six key general benchmarks using the Eleuther AI Language Model Evaluation Harness [^2] https://github.com/EleutherAI/lm-evaluation-harness, a unified framework designed to test generative language models across a vast array of evaluation tasks. For code-related tasks, we employ the BigCode Evaluation Harness [^3] https://github.com/bigcode-project/bigcode-evaluation-harness to evaluate HumanEval and MBPP, and we report the pass@1 rate of code tasks with greedy decoding.
The benchmarks used for evaluation include:
@@ 4.2 Pretrain Results
The results highlight that LLaMA Pro effectively balances natural language processing and coding capabilities. It not only preserves the general performance of its base model, LLaMA2-7B, but also surpasses it in the average performance of general language tasks. Conversely, CodeLLaMA-7B sacrifices general performance to enhance its code ability. We attribute this improvement to our expansion design, which freezes the initial LLaMA blocks to maintain their capabilities and increases the blocks to accommodate more domain-specific knowledge.
@@ 4.3 SFT Results
As seen in Figure 1, LLaMA Pro - Instruct boosts both code and math tasks to state-of-the-art performances while maintaining reliable general performance. We enhance the average performance of LLaMA2-7B-chat and CodeLLaMA-7B-instruct by 13.81% and 14.50% respectively, which highlights the benefits of balancing textual and coding abilities.
To assess the comprehensive conversational performance of the LLaMA Pro - Instruct assistant, we evaluate it using the MT-Bench with GPT-4 automatic scoring, as proposed by Vicuna Zheng et al. (2023). As depicted in Table 4, LLaMA Pro - Instruct surpasses widely used chatbots from the LLaMA community. This indicates its potential as a chatbot capable of providing helpful responses, in addition to its impressive performance in traditional benchmarks. The details of MT-Bench can be found in the Appendix C.
We use MINT-Bench  Wang et al. (2023c) to evaluate our model’s ability to solve multi-turn interactions by using tools. MINT-Bench tests LLMs’ ability to use tools by generating and executing Python code, focusing on tool-augmented task-solving and leveraging natural language feedback. MINT includes eight datasets covering reasoning, code generation, and decision-making. The details of MINT can be found in the Appendix B. The results are shown in Table 5. LLaMA Pro - Instruct achieves SOTA performance compared to similar size models in multi-turn interactions with the use of tools.
@@
Table 5: In the tool-augmented reasoning assessments, we evaluate the model’s proficiency in integrating tools into its reasoning workflow. The model’s effectiveness is measured by its success rate across various stages of interaction.
@@ 4.4 Ablation Study
Method Overall Performance (OP) Backward Transfer (BWT)
LoRA 37.1 -17.3%
SeqFT 45.5 -14.7%
Block Expansion 46.5 -14.3%
Table 6: Performance comparison of various training strategies on the TRACE benchmark following their continual learning phase with LLaMA2-7B. The table presents the Overall Performance (OP) and Backward Transfer (BWT) scores for each strategy, demonstrating the superior adaptability of the proposed block expansion training approach.
Figure 5: Training loss with varying added blocks and mixture-of-expert (MoE) expansion.
@@
Table 7: Comparison of evaluation results among several prominent code and language models. The last column represents the average of the language task average and the code task average.
@@
We evaluate various training strategies, including LoRA, fine-tuning, and the block expansion training approach that we propose, using the TRACE benchmark Wang et al. (2023b). TRACE is designed to assess continual learning in LLMs and comprises eight distinct datasets that span challenging tasks such as domain-specific tasks, multilingual capabilities, code generation, and mathematical reasoning. We assess the ability of different strategies to retain the model’s existing knowledge while incorporating new skills. Details are provided in the Appendix  13.
We employ Overall Performance (OP Chaudhry et al. (2018)) and Backward Transfer (BWT Lopez-Paz and Ranzato (2017)) scores as evaluation metrics. After incrementally learning the $t$-th task, the model’s score on the $i$-th task (where $i\leq t$) is denoted as $R^{D}_{t,i}$. The OP and BWT scores are calculated using the following formulas:
$OP_{t}=\frac{1}{t}\sum_{i=1}^{t}R_{t,i}^{D}$ (6)
$BWT_{t}=\frac{1}{t}\sum_{i=1}^{t}\left(R_{t,i}^{D}-R_{i,i}^{D}\right)$ (7)
Table 6 presents the performance of different strategies on the TRACE benchmark following their continual learning phase with LLaMA2-7B. The results show that block expansion training exhibits superior task-specific adaptability compared to sequential fine-tuning and LoRA, as evidenced by its better OP and BWT scores.
Apart from the aspect of code corpus, we explore our method on another domain: law, with the freelaw subset of Pile dataset as our pretrain corpus  Gao et al. (2020). We evaluate on UNFAIR-ToS  Lippi et al. (2019) of the LexGLUE benchmark  Chalkidis et al. (2021). The details can be found in the Appendix 14.
In our experiment, we assess the scalability of our block expansion method in terms of training loss and downstream task performance as we increase the number of added blocks. We also compare our method with the Mixture-of-Expert (MoE) expansion method Fedus et al. (2022).
We first examine the training loss with varying added blocks. As seen in Figure 5, the training loss of the models consistently decreases as training progresses, regardless of the number of added blocks. Moreover, the loss decreases more rapidly as we increase the size of the model. These findings suggest that our method exhibits strong scalability with larger models and more data. The training loss of MoE is comparable to our method with four added blocks.
However, a lower overall training loss does not necessarily guarantee superior performance on domain-specific tasks. Therefore, we evaluate models of different sizes on both general language tasks and Unfair-ToS, as shown in Table 7. All the expanded models effectively preserve the general capabilities of the initial model. For the domain-specific task, larger models achieve better performance. We find that adding eight blocks provides optimal performance with minimal cost compared to larger models, hence we adopt this as our default strategy.

### v2

@@ Abstract
Humans generally acquire new skills without compromising the old; however, the opposite holds for Large Language Models (LLMs), e.g., from LLaMA to CodeLLaMA. To this end, we propose a new post-pretraining method for LLMs with an expansion of Transformer blocks. We tune the expanded blocks using only new corpus, efficiently and effectively improving the model’s knowledge while mitigating forgetting. In this paper, we experiment on the corpus of code and math, yielding LLaMA Pro-8.3B, a versatile foundation model initialized from LLaMA2-7B, excelling in general tasks, programming, and mathematics. LLaMA Pro and its instruction-following counterpart (LLaMA Pro - Instruct) achieve advanced performance among various benchmarks, demonstrating superiority over existing open models in the LLaMA family and the immense potential of reasoning and addressing diverse tasks as an intelligent agent. Our findings provide valuable insights into integrating natural and programming languages, laying a solid foundation for developing advanced language agents that operate effectively in various environments.
LLaMA Pro: Progressive LLaMA with Block Expansion
Chengyue Wu1,2   Yukang Gan2   Yixiao Ge2thanks: Correspondence to yixiaoge@tencent.com. Zeyu Lu3   Jiahao Wang1   Ye Feng4   Ying Shan2   Ping Luo1 1The University of Hong Kong   2ARC Lab, Tencent PCG 3Shanghai Jiao Tong University   4Beijing Language and Culture University https://github.com/TencentARC/LLaMA-Pro
@@ 2 Related Work
Recent advancements in large language models have led to significant progress, with model and data scale growth driving state-of-the-art performance across various tasks Hoffmann et al. (2022); Kaplan et al. (2020); Chowdhery et al. (2023). The development of generalist models has enabled addressing diverse problems and rapid adaptation to new tasks Radford et al. (2019); Brown et al. (2020). The open-source community has further contributed by releasing powerful models like LLaMA Touvron et al. (2023) and CodeLLaMA Roziere et al. (2023). Our work builds upon these developments, providing a method for specializing LLMs in the code domain, fostering future research and applications.
@@
Language model applications typically involve a two-step process: general-domain pretraining followed by domain-specific training Roziere et al. (2023); Azerbayev et al. (2023). Fine-tuning often aims to enhance instruction-following abilities Sanh et al. (2021); Wei et al. (2021); Wang et al. (2023d) or align model outputs with human preferences Ziegler et al. (2019); Ouyang et al. (2022); Bai et al. (2022). Some research explores parameter-efficient fine-tuning methods for adapting pretrained models to new domains Houlsby et al. (2019); Hu et al. (2021); Wu et al. (2023a), while others focus on continual learning post-pretraining Wang et al. (2023b); Gupta et al. (2023); Scialom et al. (2022). Parameter-efficient tuning methods like adaptor and LoRA are generally applied during the instruction tuning phase rather than the pretraining phase. In contrast, our focus is on enhancing the capacity of LLMs by increasing their depth during continued pretraining. Our work proposes an adaptation strategy that combines continued training with general capability maintenance, allowing LLMs to specialize without sacrificing overall performance.
@@
Progressive training has gained attention for accelerating large-scale model training in computer vision Zhang et al. (2023) and NLP research Yao et al. (2023); Li et al. (2023b). Gong et al. (2019) proposed a stacking method doubling model depth successively. CompoundGrowGu et al. (2020) extends stacking with Feed-Forward Network expansion in schedule design. Shen et al. (2022) introduced a staged method supporting hidden size expansion. Bert2BERTChen et al. (2021a) and LiGO Wang et al. (2023a) accommodate all growth dimensions. Our method utilizes depth growth to maintain general performance while adapting to specific domains.
@@ 3.1 Preliminaries: The LLaMA Block
The FFN block in the LLaMA model utilizes the SwiGLU activation function, which is defined as:
@@ 3.2 Block Expansion
Given a model with blocks $(\phi_{0},\phi_{1},...,\phi_{L})$, the block expansion incorporates an identity block $\phi_{id}$ after each block in the original model, ensuring that the expanded model maintains the same output after expansion. The identity block is defined as $\phi_{id}(x)=x$ so the input and output are identical.
@@
The entire training pipeline is depicted in Figure 2. Our method concentrates on the post-pretraining stage, targeting specific domain corpora. We begin by initializing our model with large language models trained on extensive unlabeled general corpora, where all blocks will be fine-tuned. To enhance the model’s capacity for accommodating additional domain knowledge while retaining its general knowledge, we employ block expansion to increase the number of blocks in the LLM. During this process, we only fine-tune the newly added blocks while freezing the original blocks, thereby preserving the general abilities of the model.
@@ 4 Experiments
This section presents our key experimental findings. We begin with experimental settings (described in Sec. 4.1), and then verify the effectiveness of block expanded tuning after pretraining (described in Sec. 4.2). Next, we give the supervised finetuning (SFT) results (described in Sec. 3). Finally, ablation studies of the key design choices are presented (described in Sec. 4.5).
@@ 4.1 Experimental Settings
We construct a dataset that concentrates on code and math. For the code component, we rely on the Stack-dedup dataset, which is a compilation of permissively licensed source codes from GitHub. Among all the programming languages available in Stack-dedup, we specifically utilize the Python split. As for the math component, we opt for the Proof-pile-2 dataset Azerbayev et al. (2023), a 55-billion-token amalgamation of scientific papers, web data containing mathematical content, and mathematical code. The details can be found in Appendix 3.
@@
We want to highlight that our approach does not incur higher training costs, and it is worth the extra resources to achieve a better performance of the domain specific tasks in the inference.
Training stage cost: Our approach requires fewer computational resources since only the newly added blocks are tuned during training. As illustrated in Figure  4, LLaMA Pro-8B (1B parameters tuned for 80B tokens) incurs less training overhead compared to CodeLLaMA-7B (7B parameters tuned for 500B tokens). It also uses fewer resources than training domain-specific models from scratch, such as StarCoder and CrystalCoder. Despite this, our method achieves a better balance of general and domain-specific performance, offering a more cost-effective solution.
Inference stage cost: Although our method requires more resources during inference than the initial LLM, it strikes a balance between performance and efficiency. LLaMA Pro-8B outperforms larger models like LLaMA2-13B and LLaMA2-34B in the code domain while demanding significantly fewer resources during training and inference.
@@
During the instruction fine-tuning phase, we combine five data sources to create LLaMA Pro - Instruct as shown in Table 7. The final sft dataset consists of approximately 1M samples. To fine-tune the basic models, we employ specific configurations, including a batch size of 128, a sequence length of 4096, 0.03 warmup ratio, a learning rate of 2e-5, a Cosine learning rate scheduler, and bf16 mixed precision.
@@
We conduct a comparative analysis of LLaMA Pro with the latest state-of-the-art (SOTA) Large Language Models (LLMs). The evaluation is performed on six key general benchmarks using the Eleuther AI Language Model Evaluation Harness [^1] https://github.com/EleutherAI/lm-evaluation-harness, a unified framework designed to test generative language models across a vast array of evaluation tasks. For code-related tasks, we employ the BigCode Evaluation Harness [^2] https://github.com/bigcode-project/bigcode-evaluation-harness to evaluate HumanEval and MBPP, and we report the pass@1 rate of code tasks with greedy decoding. The evaluation details can be found in Appendix D.
@@ 4.2 Pretrain Results
The results highlight that LLaMA Pro effectively balances natural language processing and coding capabilities. It not only preserves the general performance of its base model, LLaMA2-7B, but also surpasses it in the average performance of general language tasks. Conversely, CodeLLaMA-7B sacrifices general performance. We attribute this improvement to our expansion design, which freezes the initial LLaMA blocks to maintain their capabilities and increases the blocks to accommodate domain-specific knowledge.
@@ 4.3 SFT Results
As seen in Figure 1, LLaMA Pro - Instruct boosts both code and math tasks to SOTA performances while maintaining reliable general performance. We enhance the average performance of LLaMA2-7B-chat and CodeLLaMA-7B-instruct by 13.81% and 14.50% respectively, which highlights the benefits of balancing textual and coding abilities.
To assess the comprehensive conversational performance of the LLaMA Pro - Instruct assistant, we evaluate it using the MT-Bench with GPT-4 automatic scoring, as proposed by Vicuna Zheng et al. (2023). As depicted in Table 2, LLaMA Pro - Instruct surpasses widely used chatbots from the LLaMA community. This indicates its potential as a chatbot capable of providing helpful responses, in addition to its impressive performance in traditional benchmarks. The details of MT-Bench can be found in the Appendix F.
We use MINT-Bench  Wang et al. (2023c) to evaluate our model’s ability to solve multi-turn interactions by using tools. MINT-Bench tests LLMs’ ability to use tools by generating and executing Python code, focusing on tool-augmented task-solving and leveraging natural language feedback. MINT includes eight datasets covering reasoning, code generation, and decision-making. The details of MINT can be found in the Appendix E. The results are shown in Table LABEL:tab:mint. LLaMA Pro - Instruct achieves SOTA performance compared to similar size models in multi-turn interactions with the use of tools.
@@
Table 3: : In the tool-augmented reasoning assessments, we evaluate the model’s proficiency in integrating tools into its reasoning workflow. The model’s effectiveness is measured by its success rate across various stages of interaction.
4.4 Mistral-Pro Results
Model ARC Hellaswag MMLU TruthfulQA Winogrande GSM8K HumanEval
Gemma-7B 61.9 82.2 64.6 44.8 79.0 50.9 32.3
Mistral-7B 60.8 83.3 62.7 42.6 78.0 39.2 28.7
Mistral-Pro (Ours) 63.2 82.6 60.6 48.3 78.9 50.6 32.9
Table 4: Comparison between the original Mistral-7B Jiang et al. (2023), Gemma-7B Team et al. (2024), and our Mistral-Pro with the Open LLM leaderboard metrics.
We experimented with block expansion on Mistral-7B  Jiang et al. (2023), training it on code and mathematics datasets. The resulting pretrained performance is detailed in Table 4, highlighting superior outcomes across various benchmarks, particularly in the domains of code and math. Notably, it demonstrates competitive results compared to the new open-source model Gemma  Team et al. (2024), while incurring significantly lower training overhead. We further utilized the MetaMath dataset  Yu et al. (2023) for supervised fine-tuning. Our approach yielded scores of 78.4 for GSM8k and 30.3 for MATH, surpassing Mistral’s scores of 77.7 and 28.2, respectively. Additional details are provided in Appendix  C.
@@ 4.5 Ablation Study
Figure 5: The training loss is analyzed with respect to the addition of varying blocks and mixture-of-expert (MoE) expansion, in conjunction with traditional training strategies such as finetuning and LoRA.
@@
[…]
Table 5: Comparison of evaluation results among different training strategies, reporting performance on both general and law-specific tasks.
Figure 6: Token distribution shift after block expansion compared to the initial LLaMA-2-7B. The proportions of unshifted, marginally shifted, and significantly shifted tokens are color-coded and presented as percentages. Frequently shifted tokens are displayed below.
@@
Apart from the aspect of code corpus, we explore our method on another domain: law, with the freelaw subset of Pile dataset as our pretrain corpus  Gao et al. (2020). We evaluate on UNFAIR-ToS  Lippi et al. (2019) of the LexGLUE benchmark  Chalkidis et al. (2021).
In our experiment, we assess the scalability of our block expansion method in terms of training loss and downstream task performance as we increase the number of added blocks. We also compare our method with the Mixture-of-Expert (MoE) expansion method Fedus et al. (2022) and traditional training strategies, such as fine-tuning and LoRA Hu et al. (2021). The details can be found in Appendix 12.
We analyze the training loss with varying added blocks (Figure 5). The loss consistently decreases during training, regardless of the number of added blocks, and decreases more rapidly with larger models. These findings indicate that our method demonstrates strong scalability with larger models and more data.
However, a lower overall training loss does not necessarily guarantee superior performance on domain-specific tasks. Therefore, we evaluate models of different sizes on both general language tasks and Unfair-ToS, as shown in Table 5. All the expanded models effectively preserve the general capabilities of the initial model. For the domain-specific task, larger models achieve better performance. We find that adding eight blocks provides optimal performance with minimal cost compared to larger models, hence we adopt this as our default strategy. The performance of MoE is comparable to our method with four added blocks. Figure 9 illustrates the differences between traditional training strategies such as fine-tuning and LoRA, and our proposed method. We observe that while LoRA effectively preserves the general ability, it struggles to model the distribution of a new domain, as also evidenced by the training loss depicted in Figure 5. In contrast, full fine-tuning results in a more significant drop in general performance. Here we use a rank of 1024 for LoRA, resulting in a number of trainable parameters comparable to our method.
In line with the approach of  Lin et al. (2023), we analyze the token distribution between the original LLaMA and LLaMA Pro to assess the similarity in their behavior when answering general questions from the Alpaca dataset Taori et al. (2023). As depicted in Figure 6, the token distribution shift between LLaMA and LLaMA Pro is subtle. Detailed information can be found in Appendix G.
@@
5 Scope and Limitations
Although our study presents a promising method for balancing general and domain-specific capabilities in LLMs, its scope is limited to the language modality, especially programming language and English. Future research could explore extending the application of our block expansion method to other domains, such as maintaining original language ability in multimodal large language modelsGe et al. (2023); Bai et al. (2023), and multi-lingual domains.